OpenStudy (jiteshmeghwal9):

What is the difference in the sum of the squares & the difference of squares of 'n' & 'a' when \[\LARGE{\frac{\sqrt{a^{n-2}.3^{a+2}}}{6^n.\left( n \over 2 \right)-\left( 1-{n \over 2} \right)}=\frac{1}{4\sqrt{2^n}}}\] & sum of 'a' & 'n' is 12:

OpenStudy (jiteshmeghwal9):

@mukushla @ParthKohli @TuringTest @amistre64 plz help:)

OpenStudy (amistre64):

that looks like a whole lotta algebra

OpenStudy (amistre64):

looks like you should prolly get rid of the radicals, turn them into exponents; cross multiply, and work it from there

OpenStudy (jiteshmeghwal9):

i need full solution plz:)

OpenStudy (amistre64):

then you really should work at it ....

OpenStudy (amistre64):

at the moment, its a bit hectic, but it can be made to look better by using basic algebraic principles

OpenStudy (jiteshmeghwal9):

@mathslover @experimentX @apoorvk can u give me full solution to this ques:)

mathslover (mathslover):

Why not i can give but only if i understand ... I am going to do that now wait

OpenStudy (jiteshmeghwal9):

k!

mathslover (mathslover):

is the answer \(\Huge{2a^2}\)

mathslover (mathslover):

@jiteshmeghwal9 ?

OpenStudy (anonymous):

what is the question? I mean , what are we supposed to find

mathslover (mathslover):

What is the difference in the sum of the squares & the difference of squares of 'n' & 'a' @vamgadu

OpenStudy (anonymous):

that means we have to find 2a^2

mathslover (mathslover):

yes hence i asked him that is the answer : 2a^2 :D

OpenStudy (anonymous):

I think we should find a more numeric answer

OpenStudy (anonymous):

:P

OpenStudy (jiteshmeghwal9):

no ! the answer is 50 :/

OpenStudy (anonymous):

ok, I'll proceed from here in this manner. the RHS is does not have a power of 3, where as the LHS has an exponent of 3. so I would say the exponent of 3 in the LHS should become 0. I am working under the assumption that a is not a multiple of 3. If I get an absurd proceeding this way I'll come back and say"Dude, my assumption was wrong. so a is a multiple of 3" Now using this assumption, power of 3 in th LHS should become 0, which means a+2-n=0 we already have a+n=12. that gives n=7 and a=5 which is consistent with the assumption that a is not a multiple of 3. so the answer become 2a^2=50

OpenStudy (anonymous):

If you still are ambiguous about the solution or you need to present it some where, I would suggest assuming a is a multiple of 3 as the first case and prove it leads to an absurd

OpenStudy (jiteshmeghwal9):

can't we find the answer without assumption

OpenStudy (anonymous):

i am not assuming any thing. I basically am dividing the solution into 2 cases. case 1: if a is a multiple of 3. proceed in this manner and this should leads to an absurd. case 2: as case 1 is an absurd, a is not a multiple of 3 and proceed in the manner I posted. I think this is the easiest way.

OpenStudy (jiteshmeghwal9):

sorry ! but i mn't getting it can u plz tell me through an easy way:)

mathslover (mathslover):

a is equal to 5 i think hence 2a^2=2*25=50

OpenStudy (jiteshmeghwal9):

ya ! it may be

OpenStudy (anonymous):

plz wait. I'll find out another way.:-)

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (jiteshmeghwal9):

@ParthKohli :)

OpenStudy (jiteshmeghwal9):

k! i have closed the ques.

mathslover (mathslover):

Why @jiteshmeghwal9 ?

mathslover (mathslover):

Is this of NTSE ?

OpenStudy (jiteshmeghwal9):

yes! @mathslover

mathslover (mathslover):

TMH /?

OpenStudy (anonymous):

dude I think there is a mistake with this question. I wrote a code in MATLAB and found the solutions are a=6.26 and n=5.74 and they are consistent with the given equation.so, check the question once and confirm. I'll retry.@jiteshmeghwal9